# Integrating Window-Based Correlated Decoding with Constant-Time Logical Gates for Large-Scale Quantum Computation

### v1

@@ Abstract
Large-scale quantum computation requires to be performed in the fault-tolerant manner. One crucial issue of fault-tolerant quantum computing (FTQC) is reducing the overhead of implementing logical gates. Recently proposed correlated decoding and “algorithmic fault tolerance” achieve fast logical gates that enables universal quantum computation. However, for circuits involving mid-circuit measurements and feedback, this approach is incompatible with window-based decoding, which is a natural requirement for handling large-scale circuits. In this letter, we propose an alternative architecture that employs delayed fixup circuits, integrating window-based correlated decoding with fast transversal gates. This design significantly reduce both the frequency and duration of correlated decoding, while maintaining support for constant-time logical gates and universality across a broad class of quantum codes. More importantly, by spatial parallelism of windows, this architecture well adapts to time-optimal FTQC, making it particularly useful for large-scale computation. Using Shor’s algorithm as an example, we explore the application of our architecture and reveals the promising potential of using fast transversal gates to perform large-scale quantum computing tasks with acceptable overhead on physical systems like ion traps.
@@
Introduction—Quantum computing offers the potential to tackle complex problems such as large integer factorization [1] and quantum simulation [2, 3]. Large-scale quantum computation relies on quantum error correction (QEC) [4, 5] and fault-tolerant quantum computing (FTQC) [6, 7]. In the conventional FTQC framework, error correction follows each logical gate to prevent error accumulation [8], typically involving $\Theta(d)$ rounds of syndrome extraction, where $d$ is the code distance. Consequently, the time of a logical gate implementation is directly relavant to $d$.
@@
Based on correlated decoding, Ref. [9] indicated that only one round of syndrome extraction is required after transversal gates, thereby achieving fast transversal gate. However, transversal gates of any QEC code are not universal [10, 11, 12]. For example, in the surface code, the implementation of logical $T$ gates requires magic states and subsequent fixup operations according to the mid-circuit measurement results [13]. Applying fast transversal gates directly to these circuits can lead to decoding failures due to insufficient syndrome extraction rounds before mid-circuit measurement.
@@
Recently, a fast $T$ gate scheme called “algorithmic fault tolerance” was proposed [14], where certain appropriate logical operators are applied during the decoding to reinterpret inconsistent measurement results of magic states. Finding these logical Pauli operators requires constructing and solving an additional linear system of equations. Particularly, the decoder and feedback process must be invoked after each mid-circuit measurement on magic states (see Fig. 1a). Since correlated decoding algorithms are typically slow, this increases the frequency and duration of decoding.
@@
More notably, this design is not compatible with the window-based decoding [15], which invokes decoder within each fixed-size window. The window-based decoding stands out as a pivotal technology, enabling the decoding of large volumes of syndrome data in a scalable and efficient manner [15, 16, 17, 18, 19]. This approach divides the continuous stream of syndrome information into manageable windows, allowing for parallel processing and reducing the complexity of decoding in large-scale quantum circuits.
@@
In this letter, we design an alternative architecture that realizes constant-time logical gates for universal quantum computing, seamlessly integrating with window-based correlated decoding. This method utilizes delayed fixup circuits to postpone the fixup until sufficient syndrome extraction rounds have occurred, thereby ensuring fault-tolerant measurements and feedback. While achieving fast logical gates and universality, the main advantage of our architecture is that it permits only a single decoder invocation for each sliding window (see Fig. 1b). In other words, syndrome data within a window is processed in a single batch, thus enhancing the efficiency of the correlated decoder in large-scale circuits.
@@
Figure 1: Comparison of correlated decoding without and with window architecture. (a) Without window architecture, correlated decoding is invoked after each mid-circuit measurement. (b) With sliding windows and delayed fixups, correlated decoding processes syndrome data within each fixed-size window.
@@
Despite decoding with temporal sliding-window being well-suited for quantum circuits with sequentially implemented logical gates, time-optimal quantum computation (TOQC), where many logical gates operate in parallel, is another common structure in large-scale quantum computing [20, 21, 22]. TOQC converts time overhead into space overhead, increasing the number of parallel qubits. Since the time for exact correlated decoding grows exponentially with the number of logical qubits [9, 16, 17], it faces significant challenges in handling TOQC circuits with large qubit counts. Interestingly, the spatial window decoding uniquely fits TOQC, which reduces the size of windows and harnesses parallelism to accelerate the decoding, thereby broadens the range of applications for window-based correlated decoding.
@@
Furthermore, we explore the application of our architecture in Shor’s algorithm as an example. Considering connectivity requirements, transversal gates are more feasible on platforms like ion traps [23, 24] or neutral atoms [25]. One QEC round on ion trap platforms is approximately 100 to 1,000 times slower than on superconducting platforms [26, 27, 28], leading to a difference in clock speed of logical gates [29]. For instance, factoring a 2048-bit integer on superconducting systems takes 8 hours [21]. One might infer that ion trap systems could require an impractical time of 800 to 8,000 hours. However, our analysis reveals that the overall runtime of the algorithm in ion-trap systems, when using fast transversal gates and TOQC circuits, is only around an order of magnitude longer than that in superconducting systems, while the space overhead does not increase significantly. Overall, we provide optimistic support for performing large-scale quantum computation tasks with acceptable space-time overhead on physical systems such as ion traps.
@@
Delayed Fixup Circuits—We illustrate the use of delayed fixup circuits for fast implementation of non-Clifford gates, specifically in surface codes, though the approach applies to other QEC codes with transversal CNOT gates and measurements [30, 31]. For surface codes, $T$ gates are implemented using the gadget in Fig. 2a, where fixup operation is classically controlled by the measurement outcome on magic state. When executing the $T$-gadget with fast transversal gates, fewer than $\Theta(d)$ syndrome extraction rounds may be insufficient to ensure fault-tolerant decoding. To accumulate $\Theta(d)$ rounds of syndrome data, computational qubits must repeatedly extract syndromes, followed by decoding and feedback, making $T$ gate implementation time potentially proportional to code distance $d$ in the worst case, rather than constant.
@@
Here, we introduce the delayed fixup circuit to achieve constant-time non-Clifford gates. The delayed fixup circuits like Fig. 2b have been elegantly designed in other previous works [20, 22, 32], whose primary purpose is to achieve gate parallelism for reducing time overhead, rather than to realize constant-time non-Clifford gates. By selecting the ancilla measurement basis, fixup decisions can be delayed. For example, in Fig. 2b, the delayed fixup circuit uses two ancilla states, |$\ket{T}=(\ket{0}+e^{i\frac{\pi}{4}}\ket{1})/{\sqrt{2}}$ and |$\ket{S}=(\ket{0}+i\ket{1})/{\sqrt{2}}$, for implementing the $T$ gate and the $S$ fixup, respectively. On the ancilla state |$\ket{S}$, the measurement is delayed until sufficient syndrome extraction rounds have been completed. Once correlated decoding provides a reliable measurement result for |$\ket{T}$, the basis of |$\ket{S}$ measurement is chosen. For circuits with multiple non-commuting $T$ gates, the |$\ket{S}$ measurement basis must be chosen sequentially. Note that on the computational qubits, only transversal CNOT gates are performed, which eliminates repeated syndrome extraction and ensures that the implementation time of the $T$ gate is irrelevant to $d$.
Figure 2: (a) Gadget for implementing $T$ gate. (b) $T$-gadget with delayed $S$ fixup. The decision for $S$ fixup is delayed until the choice of the $X$ or $Z$ measurement basis.
@@
In summary, the delayed fixup circuit ensures sufficient syndrome extraction rounds before fixup operations, making the decision on fixup fault-tolerant. Combined with transversal Clifford gates on surface codes, it is easy to confirm that this architecture achieves constant-time logical gates in a universal set. A potential concern is that keeping the |$\ket{S}$ state longer in the circuit might accumulate logical errors or increase qubit usage. However, as we will show later, with an appropriate window length, the |$\ket{S}$ state persists for at most $\Theta(d)$ rounds, preventing significant increase in logical error rates or resource overhead.
Correlated Decoding with Sliding Windows—Next, we introduce the temporal sliding windows in correlated decoding and explain its unique adaptability for delayed fixup circuits. For large-scale quantum circuits, dividing syndrome data into windows for separate processing is essential. However, applying a sliding-window architecture to fast non-Clifford gate circuits is not straightforward. The sliding window method processes entire syndrome data in each window, which is clearly incompatible with the fast $T$-gate scheme in Ref. [14], where correlated decoding is applied after each magic state measurement.

### v2

# Scalable Constant-Time Logical Gates for Large-Scale Quantum Computation Using Window-Based Correlated Decoding

@@ Abstract
Large-scale quantum computation requires to be performed in the fault-tolerant manner. One crucial challenge of fault-tolerant quantum computing (FTQC) is reducing the overhead of implementing logical gates. Recently work proposed correlated decoding and “algorithmic fault tolerance” to achieve constant-time logical gates that enables universal quantum computation. However, for circuits involving mid-circuit measurements and feedback, the previous scheme for constant-time logical gates is incompatible with window-based decoding, which is a scalable approach for handling large-scale circuits. In this work, we propose an architecture that employs delayed fixup circuits and window-based correlated decoding, realizing scalable constant-time logical gates. This design significantly reduces both the frequency and duration of decoding, while maintaining support for constant-time and universal logical gates across a broad class of quantum codes. More importantly, by spatial parallelism of windows, this architecture well adapts to time-optimal FTQC, making it particularly useful for large-scale quantum computation. Using Shor’s algorithm as an example, we explore the application of our architecture and reveals the promising potential of using constant-time logical gates to perform large-scale quantum computation with acceptable overhead on physical systems like ion traps.
I Introduction
@@
Quantum computing offers the potential to tackle complex problems such as large integer factorization [1] and quantum simulation [2, 3]. Large-scale quantum computation relies on quantum error correction (QEC) [4, 5] and fault-tolerant quantum computing (FTQC) [6, 7]. Fault-tolerant gate operations on logical qubits are typically much more complex than operations on physical qubits, consuming substantial space-time overhead. Therefore, developing low-overhead logical gates is a crucial challenge in driving FTQC toward practical implementation. In the conventional FTQC framework, error correction follows each logical gate to prevent error accumulation and propagation [8], typically involving $\Theta(d)$ rounds of syndrome extraction, where $d$ is the code distance. Consequently, the time of a logical gate implementation is directly relevant to $d$.
@@
Based on correlated decoding, Ref. [9] indicated that only one round of syndrome extraction is required after transversal gates, thereby achieving constant-time transversal gate. The fast implementation of transversal gates reduces the overall space-time overhead in fault-tolerant quantum computing (FTQC). However, transversal gates of any QEC code are not universal [10, 11, 12]. For example, for the surface code, a transversal gate set includes $\{H,S,\text{CNOT}\}$ [13], assuming a platform with all-to-all connectivity, such as trapped ions [14, 15] or neutral atoms [16]. While the implementation of logical $T$ gates requires magic states and subsequent fixup operations $S$ according to the mid-circuit measurement outcomes [17]. Directly applying constant-time transversal gates in circuits involving mid-circuit measurement and feedback is infeasible, since there are insufficient rounds of syndrome extraction to ensure reliable decoding results before making feedback decisions. The decoding result immediately after mid-circuit measurements is possibly inconsistent with subsequent decoding results obtained after sufficient syndrome extraction rounds, leading to incorrect fixup operations.
@@
Recently, a fast $T$ gate scheme called “algorithmic fault tolerance” was proposed [18], the key of which is addressing inconsistencies decoding results through classical processing. Specifically, when the measurement outcome on the magic state provided by the decoder is inconsistent with previous results, certain trivial logical Pauli operators are applied to the initial state to flip the measurement outcome on the magic state, thereby maintaining consistency. Finding these operators requires constructing and solving an additional linear system of equations during the decoding. In addition, the decoder and feedback process must be invoked after each mid-circuit measurement on magic states (see Fig. 1a). Given the execution of the correlated decoding algorithm are typically slower, this increases the frequency and duration of decoding.
@@
More notably, this design is incompatible with window-based decoding, which stands out as a pivotal technology for decoding large volumes of syndrome data in a scalable and efficient manner [19]. Window-based decoding divides the continuous stream of syndrome information into manageable windows, allowing for parallel processing to reduce the time complexity of decoding in large-scale quantum circuits [19, 20, 21, 22, 23]. In this approach, the decoder is invoked once for each fixed-size window to process the corresponding syndrome data. On the contrary, the implementation of fast $T$ gates requires invoking the decoder after each mid-circuit measurement. As window-based methods become essential for handling large-scale circuits, this inherent conflict becomes more pronounced, presenting a significant challenge to the application of previously proposed fast $T$ gate schemes.
@@
In this work, we design an architecture that realizes scalable constant-time logical gates for universal quantum computing, seamlessly integrating with window-based correlated decoding. This method utilizes delayed fixup circuits to postpone the fixup operation until sufficient syndrome extraction rounds have occurred, thereby ensuring fault-tolerant measurements and feedback in the non-Clifford gate implementation. While achieving constant-time logical gates and universality, the main advantage of our architecture is that it permits only a single decoder invocation for each sliding window (see Fig. 1b). In other words, syndrome data within a window is processed in a single batch, thus enhancing the efficiency of the correlated decoder when handling large-scale circuits.
@@
Figure 1: Comparison of correlated decoding without and with window architecture in the circuit with mid-circuit measurement and feedback. The blue blocks represent single-qubit or two-qubit logical gates and the blue dashed lines represent one round of syndrome extraction. The rectangles denote each decoder invocation and the range of syndrome data it processes. (a) Without window architecture, correlated decoding is invoked after each mid-circuit measurement. (b) In sliding window architecture, correlated decoding processes syndrome data within each fixed-size window, using delayed fixup circuits to defer the decision for feedback.
@@
By dividing syndrome data into temporal sliding windows, the correlated decoding problem can be efficiently and scalably addressed for deep quantum circuits with sequentially implemented logical gates. Nevertheless, for another widely used circuit structure in large-scale quantum computing, time-optimal quantum computation (TOQC) [24, 25, 26], the correlated decoding encounters new challenges. In TOQC, logical gates are executed in parallel to the extent possible, thereby minimizing the execution time of the circuit. The trade-off of this approach is that the time overhead is converted into space overhead, increasing the number of parallel qubits. Since the time for exact correlated decoding grows exponentially with the number of logical qubits [9, 20, 21], the decoding scale of each temporal window can easily become uncontrollable. Interestingly, we found that further dividing the windows spatially can effectively addresses this challenge. While dividing into spatial windows is not always feasible for circuits with general structures, it is particularly well-suited for TOQC. Spatial window decoding reduces the size of windows and harnesses parallelism to accelerate the decoding, thereby broadens the range of applications for window-based correlated decoding.
@@
Adapting correlated decoding to window-based approach enables its application to large-scale quantum computation in a scalable and efficient manner. Using Shor’s algorithm as an example, we explore the application of our architecture. Considering connectivity requirements, transversal gates are more feasible on platforms like ion traps [27, 28] or neutral atoms [16]. One QEC round on ion trap platforms is approximately 100 to 1,000 times slower than on superconducting platforms [29, 15, 30], leading to a difference in clock speed of logical gates [31]. For instance, factoring a 2048-bit integer on superconducting systems takes 8 hours [25]. One might infer that ion trap systems could require an impractical time of 800 to 8,000 hours. However, our analysis reveals that the overall run-time of the algorithm in ion-trap systems, when using constant-time transversal gates and TOQC circuits, is only around an order of magnitude longer than that in superconducting systems, while the space overhead does not increase significantly. Overall, we provide optimistic support for performing large-scale quantum computation tasks with acceptable space-time overhead on physical systems such as ion traps.
The remainder of this paper is structured as follows. Sec. II introduce the correlated decoding and constant-time logical gate as a foundation. Sec. III explains the proposed architecture for implementing scalable constant-time logical gates using delayed fixup circuits. Sec. IV delves into the details of temporal sliding window-based correlated decoding, explaining its mechanisms and advantages. In Sec. V, the concept of window decoding is further extended to spatial windows, with a focus on its distinctive suitability for TOQC. Sec. VI presents numerical evaluations of the performance of temporal and spatial window-based correlated decoding through two illustrative examples. Sec. VII provides a brief analysis of the impact of these results on large-scale quantum computing, using Shor’s algorithm as an example. Finally, Sec. VIII concludes and discusses this work.
II Preliminaries
This section reviews the relevant aspects of correlated decoding and explains how constant-time logical gates can be implemented within the framework of correlated decoding.
II.1 Correlated decoding
For CSS codes [6], the logical space is determined by stabilizer generators, denoted as $\{s_{x}\}$ and $\{s_{z}\}$, which represent the sets of $X$-type and $Z$-type stabilizer generators, respectively. When running multiple rounds of syndrome extraction circuits, the error syndrome $\sigma_{n,k}(s_{x/z})\in\{+1,-1\}$ is defined as the measurement outcome of the stabilizer generator $s_{x/z}$ on the $n$-th logical qubit in the $k$-th round.
Next, we assume that the logical operations in the circuit are transversal Clifford gates, followed by several rounds of syndrome extraction or transversal measurements. In the correlated decoding, the syndrome checks (or checks in brief) $C_{n,k}(s_{x/z})$, are assigned corresponding values $c_{n,k}(s_{x/z})\in\{+1,-1\}$ defined as the product of $\sigma_{n,k}(s_{x/z})$ and its previous result (if one exists), obtained by back-propagating the stabilizer operator through the circuit to the point where it was previously measured. Specifically, between the syndrome extraction in the $(k-1)$-th and $k$-th rounds, if a logical Pauli operation (or an idle operation) is applied, then
$c_{n,k}(s_{x/z})=\sigma_{n,k}(s_{x/z})\sigma_{n,k-1}(s_{x/z}).$ (1)
If logical $H_{L}$ gate is applied between the $(k-1)$-th and $k$-th rounds of syndrome extractions, then
$c_{n,k}(s_{x/z})=\sigma_{n,k}(s_{x/z})\sigma_{n,k-1}(s_{z/x}).$ (2)
If logical CNOT gate on qubit $n_{1}$ (control) and $n_{2}$ (target) is applied between the $(k-1)$-th and $k$-th rounds of syndrome extractions, then
$c_{n_{1},k}(s_{x})$ $=\sigma_{n_{1},k}(s_{x})\sigma_{n_{1},k-1}(s_{x})\sigma_{n_{2},k-1}(s_{x}),$ (3)
$c_{n_{2},k}(s_{x})$ $=\sigma_{n_{2},k}(s_{x})\sigma_{n_{2},k-1}(s_{x}),$
$c_{n_{1},k}(s_{z})$ $=\sigma_{n_{1},k}(s_{z})\sigma_{n_{1},k-1}(s_{z}),$
$c_{n_{2},k}(s_{z})$ $=\sigma_{n_{2},k}(s_{z})\sigma_{n_{2},k-1}(s_{z})\sigma_{n_{1},k-1}(s_{z}).$
It is more intuitive to explain correlated decoding on a hypergraph. In this representation, each vertex corresponds to a check $C_{i}$. Here, $C_{i}$ represents the $i$-th check in a certain sorted order of $C_{n,k}(s_{x/z})$, and its value is $c_{\textit{i}}$. Each hyperedge $e_{j}$ in the graph represents an error event, with its value $E_{j}$ being either 0 or 1, indicating whether the event does not occur or occurs, respectively. Such an error event will flip the values of the checks associated with the vertices of that hyperedge. The set of error event values $E_{j}$ that cause the flipping of the check $C_{i}$ is denoted as $I(C_{i})$. Furthermore, each hyperedge is assigned a weight $w_{j}=-\log(p_{j}/1-p_{j})$, reflecting the error probability $p_{j}$ of the error event.
As a result, the correlated decoding problem is transformed into a minimum matching problem to identify $\{E_{j}\}$ on a weighted hypergraph, as follows [9]:
maximize $\sum_{j=1}^{M}w_{j}E_{j}$ (4)
subject to $\sum_{E_{j}\in I(C_{i})}E_{j}-2K_{i}=\frac{1}{2}(1-c_{i}),$
$\qquad\qquad\quad\forall i=1,\ldots,N,$
$K_{i}\in\mathbb{Z}_{\geq 0},\quad\forall i=1,\ldots,N,$
$E_{j}\in\{0,1\},\quad\forall j=1,\ldots,M,$
where $M$ and $N$ represent the number of error events and checks, respectively, and $K_{i}$ is an integer slack variable.
According to Ref. [9], the solution to this problem can be exactly obtained using a mixed-integer programming algorithm, known as the most-likely error (MLE) decoder, which requires exponential time. Alternatively, it can be approximately solved using the hypergraph union-find (HUF) decoder [32, 33], which operates in polynomial time.
II.2 Constant-time logical gate
In the conventional FTQC framework, the decoding of each logical block is performed independently. Before executing transversal CNOT gates between logical blocks, the decoding process must first be conducted to ensure that errors are corrected before propagating through the transversal CNOT gate. Due to the imperfect measurements, this decoding requires at least $\Theta(d)$ rounds of syndrome data, where $d$ is the code distance. As a result, the time overhead for logical gates includes a factor of $\Theta(d)$ rather than being constant. It is worth mentioning that, despite differences in mechanisms, this conclusion also applies to logical operations based on lattice surgery [34, 35, 26].
However, when correlated decoding is applied to the syndrome data across multiple logical blocks, the $\Theta(d)$ rounds of syndrome extraction typically required for transversal gates are no longer necessary. In correlated decoding, transversal gates and syndrome extraction are considered as a unified process, where transversal gates are viewed as automorphic mappings on the stabilizer group of the code. By tracking these mappings, syndrome information can still be effectively extracted and used for decoding. This allows decoding of a circuit consisting of a sequence of transversal gates, where on average, only one or fewer rounds of syndrome extraction are performed per layer of transversal gates, while maintaining fault tolerance [9].
Correlated decoding enables transversal gates to be implemented fast, with a time overhead of only $\Theta(1)$. For simplicity, this paper assumes that a single round of syndrome extraction is performed immediately after each transversal gate and focuses on logical qubits encoded on surface codes, although this is not strictly necessary. A set of transversal gates for surface codes includes $\{H,S,\text{CNOT}\}$. Therefore, these logical gates can be implemented in constant time. However, this set alone is insufficient for computational universality.
Figure 2: (a) Gadget for implementing $T$ gate consuming a magic state |$\ket{T}=(\ket{0}+e^{i\frac{\pi}{4}}\ket{1})/{\sqrt{2}}$. The $S$ gate in the circuit is classically controlled by the measurement outcome of the magic state. (b) Using a trivial Pauli operator to reinterpret the measurement outcome. When twice decoding yields inconsistent results ($M_{1}$ and $M_{2}$), the measurement outcome can be reinterpreted by applying a trivial Pauli $X$ operator. The $X$ operator propagates through a CNOT gate, flipping the measurement outcome of the magic state, while leaving a $Y$ operator at the end of the circuit.
@@ II.2 Constant-time logical gate
To implement constant-time logical gates in a universal gate set, it is beneficial to consider the $T$ gate implementation using the gadget shown in Fig. 2a. Note that the fixup $S$ operation here is classically controlled by the measurement outcome on the magic state |$\ket{T}$, which depends on decoding the syndrome data prior to the mid-circuit measurement. The challenge is that if the $T$-gadget is executed with only one round of syndrome extraction per transversal gate, the syndrome data may be insufficient to ensure fault-tolerant decoding. A naive solution is to accumulate $\Theta(d)$ rounds of syndrome data by repeatedly extracting syndromes before performing decoding and feedback on the computational qubits. However, this approach potentially makes the implementation time of the $T$ gate proportional to the code distance $d$ in the worst case, rather than constant.
Ref. [18] proposes a constant-time $T$ gate scheme, illustrated with the example shown in Fig.2b. Maintaining one round of syndrome extraction per transversal gate, if the outcome of the mid-circuit measurement is later proven incorrect by subsequent decoding, a trivial $X_{L}$ operator is applied to the first logical qubit. This operator flips the measurement outcome and leaves only a Pauli $Y_{L}$ operator on the first logical qubit. This ensures that the earlier decoding inaccuracies can be effectively corrected in subsequent steps. For more general circuits with input stabilizer states, finding such logical Pauli operators requires solving a system of linear equations. Ref. [18] demonstrates that the probability of such a system being unsolvable decreases exponentially with increasing code distance, thereby ensuring fault tolerance.
Remarkably, in this scheme, the decoding must be performed immediately after each mid-circuit measurement, which makes it challenging to process large amounts of syndrome data using a manageable window-based approach. First, due to the uncertainty in mid-circuit measurement positions, a fixed-step sliding window no longer adapts effectively. Secondly, if a decoding window of size $\Theta(d)$ is used for each mid-circuit measurement, these windows will have significant overlapping regions. As a result, the same syndrome data will be processed repeatedly in multiple windows, reducing decoding efficiency.
III Constant-time non-Clifford gates with delayed fixup circuit
Here we propose an scalable constant-time implementation of non-Clifford gates using a delayed fixup circuit. We focus on surface codes to explain this approach, though it applies to other QEC codes that promise transversal CNOT gate and logical $X$ and $Z$ measurements [36, 37].
When implementing non-Clifford gates such as $T$ gates or Toffoli gates on surface-code logical qubits, it is typically required to perform fixup operations classically controlled by measurements on magic states. By selecting the ancilla measurement basis, fixup decisions can be delayed. For example, in Fig. 3a, the delayed fixup circuit uses two ancilla states, |$\ket{T}=(\ket{0}+e^{i\frac{\pi}{4}}\ket{1})/{\sqrt{2}}$ and |$\ket{S}=(\ket{0}+i\ket{1})/{\sqrt{2}}$, for implementing the $T$ gate and the $S$ fixup operation, respectively. If a $Z$ measurement is performed on the |$\ket{S}$ state, the $S$ fixup operation is applied (up to a Pauli $Z$ operation). Conversely, if the $S$ fixup operation is not needed, an $X$ measurement is performed on the |$\ket{S}$ state. Since the timing of the measurement can be arbitrarily delayed, the decision to apply the fixup operation can also be postponed until it is actually required. For circuits with multiple non-commuting $T$ gates, the |$\ket{S}$ measurement basis must be chosen sequentially, as illustrated in Fig. 3b.
Note that delayed fixup circuits are not introduced for the first time in this work. They have been elegantly designed in several previous literatures [24, 26, 38] and have been applied in various FTQC-related works [24, 25]. However, their primary purpose in these contexts has been to enable gate parallelism for reducing time overhead, rather than to realize constant-time non-Clifford gates. These results allows us to easily utilize well-designed delayed fixup circuits to implement common non-Clifford gates in constant time.
@@ III Constant-time non-Clifford gates with delayed fixup circuit
To achieve constant-time logical gates, single rounds of syndrome extraction are performed after each transversal CNOT gate on both the computational qubits and the ancilla state |$\ket{T}$. On the ancilla state |$\ket{S}$, the measurement is delayed until sufficient syndrome extraction rounds have been completed. Once correlated decoding provides a reliable measurement outcome for |$\ket{T}$, the basis of |$\ket{S}$ measurement can be chosen (as detailed in the next section). Overall, on the computational qubits, transversal CNOT gates consume constant time overhead and $Z$ fixup can be achieved in the Pauli frame [39, 40], which ensures that the implementation time of the logical $T$ gate is irrelevant to $d$.
@@
A potential concern is that keeping the |$\ket{S}$ state longer in the circuit might accumulate logical errors or increase time of qubit occupancy. However, as we will explain in the window-based decoding, with an appropriate window length, the |$\ket{S}$ state persists for at most $\Theta(d)$ rounds, preventing significant increase in logical error rates or resource overhead.
Figure 3: (a) $T$-gadget with delayed $S$ fixup. The decision for $S$ fixup is delayed until the choice of the $X$ or $Z$ measurement basis. (b) When multiple non-commuting $T$ gates are applied sequentially, the measurement basis for the |$\ket{S}$ state must be determined sequentially. The $T$-gadget is abstracted to represent, showing only the ancilla |$\ket{S}$ state while omitting the |$\ket{T}$ state.
In summary, the delayed fixup circuit ensures sufficient syndrome extraction rounds before fixup operations, making the decision on fixup fault-tolerant. More importantly, this architecture is compatible with window-based decoding methods (see the next section), thus addressing the scalability issue. Combined with constant-time transversal Clifford gates on surface codes [9], it is easy to confirm that this architecture achieves constant-time logical gates in a universal gate set. In addition, although here we focuses on the implementation of constant-time logical $T$ gates, generalizing the approach is straightforward. By utilizing the delayed fixup circuits proposed in Ref.[38] and Ref.[24], logical Toffoli gates and other non-Clifford gates can also be realized within constant time using this architecture.
IV Correlated Decoding with Sliding Windows
Now we introduce the temporal sliding windows in correlated decoding and explain its unique adaptability for the design in the previous section. To be precise, the window refers to a subregion composed of syndrome checks, which are arranged in temporal and spatial order within a three-dimensional space-time diagram. Hereafter, assuming the output of correlated decoding is $\{E_{j}\}$, we refer to the hyperedges $e_{j}$ satisfying $E_{j}=1$ as the decoding results for each windows.